If

P1
and

P2
are the pressure above and below the wing of an aeroplane.

ρ
is the density of air,

A
is the area of the wing and

v1
and

v2
are the speeds of air above and below the wing. Then dynamic lift of aeroplane is
1
Expert's answer
2021-07-13T13:26:35-0400

Given, Pressure - P1, P2

Speed - V1, V2

Area - A

Bernoulli's Law


If P2 > P1 & V1 > V2


F = P1"\rho"gh + "\frac{1}{2}""\rho"V12


P1"\rho"gh1 + "\frac{1}{2}""\rho"V12 = P2"\rho"gh2 + "\frac{1}{2}""\rho"V22


P2 - P1 = "\frac{1}{2}""\rho"V12 - "\frac{1}{2}""\rho"V22


A"\Delta"P = A[ "\frac{1}{2}""\rho"(V12 - V22 ]


F = AP = A[ "\frac{1}{2}""\rho"(V12 - V22 ]


Hence, A[ "\frac{1}{2}""\rho"(V12 - V22 ] will be the dynamic lift of airplane.